OpenStudy (anonymous):

Solve : 1/((1/s)+(10/s)) please show all work ...kind regards

OpenStudy (jdoe0001):

if I had say 2 fractions, what would be my LCD \(\bf \cfrac{3}{4}, \cfrac{5}{8} \implies \cfrac{}{lcd?}\)

OpenStudy (anonymous):

32

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{s}+\cfrac{10}{s} \implies \cfrac{}{lcd?}\)

OpenStudy (anonymous):

s

OpenStudy (jdoe0001):

sorry for the spacing out :( yes the lcd is "s"

OpenStudy (jdoe0001):

\(\bf \large \cfrac{1}{\frac{1}{s}+\frac{10}{s}}\\ \color{blue}{\cfrac{1}{s}+\cfrac{10}{s} \implies \cfrac{1+10}{s} \implies \cfrac{11}{s}}\\ \large \cfrac{1}{\frac{1}{s}+\frac{10}{s}} \implies \cfrac{1}{\frac{11}{s}} \implies \cfrac{1}{1} \times\cfrac{s}{11}\)

OpenStudy (anonymous):

excellent....thank you jdoe kind regards

OpenStudy (jdoe0001):

yw